40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 748     Entry time: Mon Jul 28 15:54:04 2008     In reply to: 738     Reply to this: 776
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ More on the beat / the PLL setup 
Alberto and Koji,

Last Friday evening, Koji found that the power adj setting (indicated by ADJ) of the NPRO was somehow set to be 
ADJ=-45 and yielded the output power of about 200mW instead of 700mW. This is not good because too small pump power 
varies thermal conditions of the crystal such as thermal lensing, thermal gradient, and os on. The ADJ setting and the 
crystal temperature had been restored to ADJ=0 and LT=~48deg (nominal of the controller), respectively.

Today we tried the quest of the beating again and the above power setting helped a lot! The beating was immediately 
found at LT=48.55deg that is very close to the laser's nominal temp. Also the beating got significantly bigger. 
After the alignment adjustment 50%-intenisity modulated signal was obtained. From the power calculation it was 
estimated that the power coupling of the injected beam is to be 12%~13%. This not so good yet, but something which we 
can work.

This time the modulation structure of the PSL beam was clearly observed. I could obtain the beating of the injection 
beam with the carrier, the upper/lower sidebands of the 33MHz and 166MHz modulations, and the 2nd order of the 
33MHz. They were beautiful as if working with an OSA. Very nice.

In reality, those additional intenisty modulations as well as the residual 33MHz signal from the main IFO are 
disturbing for the PLL to be locked at the proper frequency. So, now Alberto is working on a passive LPF with 
notch at 33MHz. The design was already done. This allows us to work up to 20MHz and at the same time, provides 
60dB attenuation at 33MHz (in principle). Very cool.

Koji, on the other hand, continued to work with the PLL servo with some ready-made passive filters. Owing to the 
fillters, the error signal was cleaner and the PLL was locked at the proper frequency. The PLL setup is as attatched. 
Sideband rejection filter will be replaced to Alberto's one. The photo is the display of the RF spectrum analyzer with 
beat locked at 8MHz.

So the next step, we try to find the resonances of the arm cavity with the injection beam once the IFO comes back.

At the last of the experiment "Last autoalignment" was restored, the flipper for the 
inj beam was down, and the shutter for the NPRO was closed.
Attachment 1: PLL_setup.png  39 kB  Uploaded Mon Jul 28 18:39:05 2008  | Show | Hide all | Show all
Attachment 2: beat_at_8MHz.jpg  78 kB  Uploaded Mon Jul 28 18:39:13 2008  | Hide | Hide all | Show all
beat_at_8MHz.jpg
ELOG V3.1.3-